OpenStudy (anonymous):

put in exponetial form 5√6:(5^(4))

OpenStudy (anonymous):

dont change it this time!

OpenStudy (anonymous):

\[5\sqrt[6]{5^{4}}\]

OpenStudy (anonymous):

5^(1+(4/6)) = 5^(10/6)

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

simplest radical form cannot have common factor in index and radicand.

OpenStudy (anonymous):

divide both by 3 to get \[5\sqrt[3]{5^2}\]